OpenStudy (anonymous):

Evaluate the expression below when y = -3 2y+7/-4y - 10 A) -1/22 B) -13/2 C) 1/2 D) 1

OpenStudy (anonymous):

Just plug in -3 for y to get your answer

OpenStudy (anonymous):

2(-3)+7=1 -4(-3)-10=2 1/2 is your answer

OpenStudy (anonymous):

Thank you very much

OpenStudy (anonymous):

Your welcome, anytime :D